# Contrastive Learning on Multimodal Analysis of Electronic Health Records

### v1

@@ 2.4 Comparison between CLAIME and Simple Concatenation
We next contrast CLAIME with the simple approach of ignoring between-modality differences between different modalities. Dealing with multimodal data often presents difficulties, leading to conventional methods that overlook the differences between various modalities. A basic strategy commonly adopted is to simply merge the two modalities through direct concatenation, after which algorithms initially intended for unimodal data are applied. However, this rudimentary treatment of multimodal data may lead to substantial bias due to the inherent heterogeneity between different modalities. To illustrate, consider the concatenated data for the $i$-th patient represented as
$\{w_{i,t}\}_{t\in[T_{i}]}=\Big{(}w_{i,1}^{(1)},\dots,w_{i,T_{i}^{(1)}}^{(1)},w_{i,1}^{(2)},\dots,w_{i,T_{i}^{(2)}}^{(2)}\Big{)},\text{ where }T_{i}=T_{i}^{(1)}+T_{i}^{(2)}\,.$
@@
A popular method to handle such data is the SVD-PMI algorithm, as referenced in (Levy and Goldberg,, 2014; Gan et al.,, 2023). In this context, we establish the co-occurrence matrices for the concatenated dataset as follows:
$\mathbf{C}=\begin{bmatrix}\mathbf{C}^{(1,1)}&\mathbf{D}^{(1,2)}\\ \mathbf{D}^{(2,1)}&\mathbf{C}^{(2,2)}\end{bmatrix},$
where $\mathbf{C}^{(M,M)}$ and $\mathbf{D}^{(M,M^{\prime})}$ are defined in Section 2.3. Subsequently, the empirical concatenated PMI matrix, denoted as $\widehat{\mathbb{PMI}}=\{\widehat{\mathbb{PMI}}(w,w^{\prime})\}_{w,w^{\prime}\in[d]}$, is formulated as
$\widehat{\mathbb{PMI}}(w,w^{\prime})$ $=\log\frac{\mathbf{C}(w,w^{\prime})\mathbf{C}(\cdot,\cdot)}{\mathbf{C}(w,\cdot)\mathbf{C}(w^{\prime},\cdot)},$ (2.5)
@@
where $\mathbf{C}(w,\cdot)=\sum_{w^{\prime}=1}^{d}\mathbf{C}(w,w^{\prime})$ and $\mathbf{C}(\cdot,\cdot)=\sum_{w=1}^{d}\mathbf{C}(w,\cdot)$. Following this, we conduct a rank-$p$ eigen-decomposition of $\widehat{\mathbb{PMI}}$, represented as $\mathbf{\widehat{U}}_{\textnormal{Con}}\widehat{\bm{\Lambda}}_{\textnormal{Con}}\mathbf{\widehat{U}}_{\textnormal{Con}}$. The estimator of $\mathbf{V}^{\star}$ is then achieved by setting $\mathbf{\widehat{V}}_{\textnormal{Con}}=\mathbf{\widehat{U}}_{\textnormal{Con}}\widehat{\bm{\Lambda}}_{\textnormal{Con}}^{1/2}$. We refer to this method as “Concate”.
@@
$\quad\quad+\frac{1}{\sum_{i=1}^{n}\sum_{j\neq i}^{n}T_{i}T_{j}}\sum_{i\neq j}\sum_{t\in[T_{i}]}\sum_{s\in[T_{j}]}\frac{\langle\mathbf{v}_{w_{i,t}},\mathbf{v}_{w_{j,s}}\rangle}{\gamma_{w_{i,t}}\gamma_{w_{j,s}}}+\frac{\lambda}{2}\|\mathbf{V}\mathbf{V}^{\top}\|_{\mathrm{F}}^{2},$
@@
$\mathbf{D}=\begin{bmatrix}\mathbf{C}^{(1)}&\mathbf{C}^{(c)}\\ \mathbf{C}^{(c)\top}&\mathbf{C}^{(2)}\end{bmatrix}\text{ with }\mathbf{C}^{(M)}(w,w^{\prime})=\sum_{i=1}^{n}\sum_{j\neq i}^{n}\mathbf{C}^{(M,M)}_{i,j}(w,w^{\prime})\text{ for }w,w^{\prime}\in{\cal W}^{(M)}.$
@@
We have
$\mathcal{L}_{\textnormal{CL}}(\mathbf{V})$ $=\frac{\lambda}{2}\Big{\|}\mathbf{V}\mathbf{V}^{\top}-\frac{1}{\lambda}\widehat{\mathbb{PMI}}_{\textnormal{CL}}\Big{\|}_{\mathrm{F}}^{2}+(\textnormal{constant}),$
where $\widehat{\mathbb{PMI}}_{\textnormal{CL}}=\{\widehat{\mathbb{PMI}}_{\textnormal{CL}}(w,w^{\prime})\}_{w,w^{\prime}\in[d]}$ with
$\widehat{\mathbb{PMI}}_{\textnormal{CL}}(w,w^{\prime}):=\frac{\mathbf{C}(\cdot,\cdot)\mathbf{C}(\cdot,\cdot)}{\mathbf{C}(w,\cdot)\mathbf{C}(w^{\prime},\cdot)}\quantity(\frac{\mathbf{C}(w,w^{\prime})}{\mathbf{C}(\cdot,\cdot)}-\frac{\mathbf{D}(w,w^{\prime})}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{1}^{(1,2)})}).$
@@ 3 Theoretical Analysis
In this section, we investigate the theoretical properties of CLAIME and draw comparisons with the Concate and CL methods. First, let us establish some fundamental concepts. For $w\in{\cal W}^{(M)}$, and $t\in[T_{i}^{(M)}]$, denote $X_{i,w}^{(M)}(t)=\mathbb{I}\{w^{(M)}_{i,t}=w\}$. According to the model (2.1), conditioned on $\bm{c}_{i}$ and ${\bm{\epsilon}}_{i}^{(M)}$, the variable $X_{i,w}^{(M)}(t)$ follows a Bernoulli distribution, with the probability of occurrence given by
@@
$p_{i,w}^{(M)}=\mathbb{E}[X_{i,w}^{(M)}(t)|\bm{c}_{i},{\bm{\epsilon}}_{i}^{(M)}]=\frac{\exp(\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w})}{\sum_{w^{\prime}\in{\cal W}^{(M)}}\exp(\langle\mathbf{v}_{w^{\prime}}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w^{\prime}})}\,.$ (3.1)
Here, the probability $p_{i,w}^{(M)}$ is a function of the discourse vector $\bm{c}_{i}$ and the noise term ${\bm{\epsilon}}_{i}^{(M)}$, and is independent of $t$. Next, we denote the expected value of $p_{i,w}^{(M)}$ as $p_{w}^{(M)}=\mathbb{E}[p_{i,w}^{(M)}]$. Further, we define the co-occurrence of a feature $w\in{\cal W}^{(M)}$ at time $t$ for the $i$-th patient and feature $w^{\prime}\in{\cal W}^{(M^{\prime})}$ at time $s$ for the $j$-th patient as $X_{i,j,w,w^{\prime}}^{(M,M^{\prime})}(t,s)=\mathbb{I}\{w^{(M)}_{i,t}=w,w^{(M^{\prime})}_{j,s}=w^{\prime}\}$ for $t\neq s$. This variable follows a Bernoulli distribution, with the probability of occurrence denoted as $p_{i,j,w,w^{\prime}}^{(M,M^{\prime})}$. Furthermore, we define the expected value for the case of $i=j$ as $p_{w,w^{\prime}}^{(M,M^{\prime})}=\mathbb{E}[p_{i,i,w,w^{\prime}}^{(M,M^{\prime})}]$. It is important to note that when $i\neq j$, the expected value $\mathbb{E}[p_{i,j,w,w^{\prime}}^{(M,M^{\prime})}]$ equals the product $p_{w}^{(M)}p_{w^{\prime}}^{(M^{\prime})}$. Now, we introduce the population PMI matrix $\mathbb{PMI}$ defined as
$\mathbb{PMI}:=\begin{pmatrix}\mathbb{PMI}^{(1,1)}&\mathbb{PMI}^{(1,2)}\\ \mathbb{PMI}^{(2,1)}&\mathbb{PMI}^{(2,2)}\end{pmatrix}\text{ with }\mathbb{PMI}^{(M,M^{\prime})}(w,w^{\prime})=\log\frac{p_{w,w^{\prime}}^{(M,M^{\prime})}}{p_{w}^{(M)}p_{w^{\prime}}^{(M^{\prime})}}$ (3.4)
@@
for $w\in{\cal W}^{(M)}$, $w^{\prime}\in{\cal W}^{(M^{\prime})}$ with $M,M^{\prime}\in\{1,2\}$. Before proceeding with our theoretical analysis, let us establish some general assumptions. Based on our data generation model, we can only identify the left singular space and the singular values of $\mathbf{V}_{1}^{\star}$ and $\mathbf{V}_{2}^{\star}$. Thus, without loss of generality, we can assume that the singular value decomposition of $\mathbf{V}_{M}^{\star}$ is $\mathbf{V}_{M}^{\star}=\mathbf{U}_{M}^{\star}\bm{\Lambda}_{M}^{\star}$ for $M\in\{1,2\}$, where $\mathbf{U}_{M}^{\star}\in\mathcal{O}_{d_{M},p}$ represents the left singular vectors, and $\bm{\Lambda}_{M}^{\star}\in\mathbb{R}^{p\times p}$ is a diagonal matrix containing the corresponding singular values. The embedding can only be identified up to a mean shift. For simplicity, we assume $\mathbf{V}_{M}^{\star\top}\bm{1}_{d_{M}}=\bm{0}$.
@@
Assumption 3.1 is widely known as the incoherence constant condition in the literature appearing in matrix completion (Candès and Recht,, 2009), PCA (Zhang et al.,, 2022), and the analysis of representation learning algorithm (Ji et al.,, 2021).
We introduce the notation $\underline{d}=\min(d_{1},d_{2})$, and ${\bm{\Sigma}}=\mathrm{diag}({\bm{\Sigma}}_{1},{\bm{\Sigma}}_{2})$.
@@
Assume that $\log d\ll d_{1}\wedge d_{2}$, $p\log^{12}d\ll d_{1}\wedge d_{2}$, $p\ll(\log\log d)^{2}$, $n\gg p^{2}d^{5}\log^{2}d$ and $p^{3}\log^{6}d\ll\underline{d}^{3/2}/d^{1/2}$.
Assumption 3.2 imposes a condition on the dimension of the low-rank representation. This assumption is technically necessary in bounding the normalizing constant of the log-linear word production model (Lemma S5.2). Although, we can relax the assumption of $p\ll(\log\log d)^{2}$ into $p\ll\log^{2}d$, we stick to the current assumption for the sake of clarity.
@@
Assumption 3.3 implies that the signal-to-noise ratio is bounded above and below. Note that the signal, measured by the smallest positive singular value of $(1/p)\mathbf{V}_{M}^{\star}\mathbf{V}_{M^{\prime}}^{\star\top}$ for $M,M^{\prime}\in\{1,2\}$ is of order $1/p$ from Assumption 3.1.
@@
$\Big{\|}\mathbb{PMI}^{(M,M)}-\quantity(\frac{1}{p}\mathbf{V}_{M}^{\star}\mathbf{V}_{M}^{\star\top}+{\bm{\Sigma}}_{M})\Big{\|}$ $\lesssim\frac{p^{3/2}}{d_{M}}\log^{6}d,\ \ M\in\{1,2\}$
$\Big{\|}\mathbb{PMI}^{(1,2)}-\frac{1}{p}\mathbf{V}_{1}^{\star}\mathbf{V}_{2}^{\star\top}\Big{\|}$ $\lesssim\frac{p^{3/2}}{\underline{d}}\log^{6}d,$
and hence
$\Big{\|}\mathbb{PMI}-\quantity(\frac{1}{p}\mathbf{V}^{\star}\mathbf{V}^{\star\top}+{\bm{\Sigma}})\Big{\|}$ $\lesssim\frac{p^{2}}{\underline{d}}\log^{6}d.$
The proof of Theorem 3.1 is presented in Supplementary S4.1. This theorem demonstrates that the PMI matrix for each modality can be closely approximated by the embedding space plus the noise covariance matrix ${\bm{\Sigma}}_{M}$. When the noise level ${\bm{\Sigma}}_{M}$ is substantial, the PMI value $\mathbb{PMI}^{(M,M)}(w,w^{\prime})$ for a single modality may significantly deviate from $\langle\mathbf{v}_{w}^{\star},\mathbf{v}_{w^{\prime}}^{\star}\rangle/p$. On the other hand, the cross-modal PMI matrix $\mathbb{PMI}^{(1,2)}$ can be directly approximated by $\mathbf{V}_{1}^{\star}\mathbf{V}_{2}^{\star\top}/p$. The theorem below affirms that $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$ is an effective estimator for $\mathbb{PMI}^{(1,2)}$.
@@
Suppose that $n\gg d^{2}\log^{2}d$. Under Assumptions 3.1-3.3 and 3.5, we have
$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}-\mathbb{PMI}^{(1,2)}}_{\max}$ $\lesssim\frac{d_{1}d_{2}}{\sqrt{n}}\log d+\frac{p^{3/2}}{\underline{d}^{3/2}}\log^{6}d\ \ \text{ and }$
@@
$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}-\mathbb{PMI}^{(1,2)}}$ $\lesssim\frac{\underline{d}^{3/2}d}{\sqrt{n}}\log d+\frac{p^{3/2}}{\underline{d}}\log^{6}d$
with probability $1-\exp(-\Omega(\log^{2}d))$.

### v2

@@ 2.3 Comparison between CLAIME and Baseline Approaches
We next contrast CLAIME with the baseline approaches of ignoring between-modality differences. Dealing with multimodal data often presents difficulties, leading to conventional methods that overlook the differences between various modalities. A basic strategy commonly adopted is to simply merge the two modalities through direct concatenation, after which algorithms initially intended for unimodal data are applied. However, this rudimentary treatment of multimodal data may lead to substantial bias due to the inherent heterogeneity between different modalities.
@@
To illustrate, consider the concatenated data for the $i$-th patient represented as $\{w_{i,t}\}_{t\in[T_{i}]}=\Big{(}w_{i,1}^{(1)},\dots,w_{i,T_{i}^{(1)}}^{(1)},w_{i,1}^{(2)},\dots,w_{i,T_{i}^{(2)}}^{(2)}\Big{)},\text{ where }T_{i}=T_{i}^{(1)}+T_{i}^{(2)}\,.$ A popular method to handle such data is the SVD-PMI algorithm, as referenced in Levy and Goldberg, (2014); Gan et al., (2025). In this context, we establish the co-occurrence matrices for the concatenated dataset as follows:
$\mathbf{C}=\begin{bmatrix}\mathbf{C}^{(1,1)}&\mathbf{D}^{(1,2)}\\ \mathbf{D}^{(2,1)}&\mathbf{C}^{(2,2)}\end{bmatrix}$.
@@
Subsequently, the empirical concatenated PMI matrix, denoted as $\widehat{\mathbb{PMI}}=\{\widehat{\mathbb{PMI}}(w,w^{\prime})\}_{w,w^{\prime}\in[d]}$, is formulated as $\widehat{\mathbb{PMI}}(w,w^{\prime})=\log\frac{\mathbf{C}(w,w^{\prime})\mathbf{C}(\cdot,\cdot)}{\mathbf{C}(w,\cdot)\mathbf{C}(w^{\prime},\cdot)}$, where $\mathbf{C}(w,\cdot)=\sum_{w^{\prime}=1}^{d}\mathbf{C}(w,w^{\prime})$ and $\mathbf{C}(\cdot,\cdot)=\sum_{w=1}^{d}\mathbf{C}(w,\cdot)$. Following this, we conduct a rank-$p$ eigen-decomposition of $\widehat{\mathbb{PMI}}$, represented as $\mathbf{\widehat{U}}_{\textnormal{Con}}\widehat{\bm{\Lambda}}_{\textnormal{Con}}\mathbf{\widehat{U}}_{\textnormal{Con}}^{{\scriptscriptstyle\sf T}}$. The estimator of $\mathbf{V}^{\star}$ is then achieved by setting $\mathbf{\widehat{V}}_{\textnormal{Con}}=\mathbf{\widehat{U}}_{\textnormal{Con}}\widehat{\bm{\Lambda}}_{\textnormal{Con}}^{1/2}$. We refer to this method as “Concate”.
@@
$\quad\quad+\frac{1}{\sum_{i=1}^{n}\sum_{j:j\neq i}^{n}T_{i}T_{j}}\sum_{i,j:i\neq j}\sum_{t\in[T_{i}]}\sum_{s\in[T_{j}]}\frac{\langle\mathbf{v}_{w_{i,t}},\mathbf{v}_{w_{j,s}}\rangle}{\gamma_{w_{i,t}}\gamma_{w_{j,s}}}+\frac{\lambda}{2}\|\mathbf{V}\mathbf{V}^{\scriptscriptstyle\sf T}\|_{\mathrm{F}}^{2},$
@@
$\mathbf{D}=\begin{bmatrix}\mathbf{C}^{(1)}&\mathbf{C}^{(c)}\\ \mathbf{C}^{(c){\scriptscriptstyle\sf T}}&\mathbf{C}^{(2)}\end{bmatrix}\text{ with }\mathbf{C}^{(M)}(w,w^{\prime})=\sum_{i=1}^{n}\sum_{j:j\neq i}^{n}\mathbf{C}^{(M,M)}_{i,j}(w,w^{\prime})\text{ for }w,w^{\prime}\in{\cal W}^{(M)}.$
@@
We have $\mathcal{L}_{\textnormal{CL}}(\mathbf{V})=\frac{\lambda}{2}\Big{\|}\mathbf{V}\mathbf{V}^{\scriptscriptstyle\sf T}-\frac{1}{\lambda}\widehat{\mathbb{PMI}}_{\textnormal{CL}}\Big{\|}_{\mathrm{F}}^{2}+(\textnormal{constant})$, where $\widehat{\mathbb{PMI}}_{\textnormal{CL}}=\{\widehat{\mathbb{PMI}}_{\textnormal{CL}}(w,w^{\prime})\}_{w,w^{\prime}\in[d]}$ with
$\widehat{\mathbb{PMI}}_{\textnormal{CL}}(w,w^{\prime}):=\frac{\mathbf{C}(\cdot,\cdot)\mathbf{C}(\cdot,\cdot)}{\mathbf{C}(w,\cdot)\mathbf{C}(w^{\prime},\cdot)}\quantity(\frac{\mathbf{C}(w,w^{\prime})}{\mathbf{C}(\cdot,\cdot)}-\frac{\mathbf{D}(w,w^{\prime})}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{2}^{(1,2)2})}).$
@@ 3 Theoretical Analysis
In this section, we establish the theoretical properties of the proposed CLAIME algorithm and provide a comparative analysis with the two representative baselines discussed in Section 2.3: the simple concatenation approach (Concate) and conventional contrastive learning (CL). Our goal is to articulate the statistical principles that underlie CLAIME, thereby clarifying both why and how it achieves performance gains relative to these alternatives.
3.1 Theoretical Properties of CLAIME
First, we introduce some basic definitions. For $w\in{\cal W}^{(M)}$ and $t\in[T_{i}^{(M)}]$, define $X_{i,w}^{(M)}(t)=\mathbb{I}\{w^{(M)}_{i,t}=w\}$. According to model (2.1), conditioned on $\bm{c}_{i}$ and ${\bm{\epsilon}}_{i}^{(M)}=(\epsilon_{i,w}^{(M)})_{w\in{\cal W}^{(M)}}$, the variable $X_{i,w}^{(M)}(t)$ follows a Bernoulli distribution with mean
@@ 3.1 Theoretical Properties of CLAIME
$p_{i,w}^{(M)}=\mathbb{E}[X_{i,w}^{(M)}(t)\mid\bm{c}_{i},{\bm{\epsilon}}_{i}^{(M)}]=\frac{\exp(\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w}^{(M)})}{\sum_{w^{\prime}\in{\cal W}^{(M)}}\exp(\langle\mathbf{v}_{w^{\prime}}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w^{\prime}}^{(M)})}.$ (3.1)
The probability $p_{i,w}^{(M)}$ depends on the discourse vector $\bm{c}_{i}$ and the noise vector ${\bm{\epsilon}}_{i}^{(M)}$, and is independent of $t$. Let $p_{w}^{(M)}:=\mathbb{E}[p_{i,w}^{(M)}]$.
Next, define the co-occurrence of $w\in{\cal W}^{(M)}$ at time $t$ for patient $i$ and $w^{\prime}\in{\cal W}^{(M^{\prime})}$ at time $s$ for patient $j$ as $X_{i,j,w,w^{\prime}}^{(M,M^{\prime})}(t,s)=\mathbb{I}\{w^{(M)}_{i,t}=w,w^{(M^{\prime})}_{j,s}=w^{\prime}\}$ for $t\neq s$, with mean $p_{i,j,w,w^{\prime}}^{(M,M^{\prime})}$. When $i=j$, define $p_{w,w^{\prime}}^{(M,M^{\prime})}:=\mathbb{E}[p_{i,i,w,w^{\prime}}^{(M,M^{\prime})}]$. For $i\neq j$, we have $\mathbb{E}[p_{i,j,w,w^{\prime}}^{(M,M^{\prime})}]=p_{w}^{(M)}p_{w^{\prime}}^{(M^{\prime})}$. We then define the population PMI matrix $\mathbb{PMI}$ as
$\mathbb{PMI}:=\begin{pmatrix}\mathbb{PMI}^{(1,1)}&\mathbb{PMI}^{(1,2)}\\ \mathbb{PMI}^{(2,1)}&\mathbb{PMI}^{(2,2)}\end{pmatrix},\quad\mathbb{PMI}^{(M,M^{\prime})}(w,w^{\prime})=\log\frac{p_{w,w^{\prime}}^{(M,M^{\prime})}}{p_{w}^{(M)}p_{w^{\prime}}^{(M^{\prime})}}$ (3.2)
for $w\in{\cal W}^{(M)}$, $w^{\prime}\in{\cal W}^{(M^{\prime})}$, and $M,M^{\prime}\in\{1,2\}$.
@@
Before proceeding, we clarify our identifiability assumptions. From (2.1), only the left singular space and singular values of $\mathbf{V}_{1}^{\star}$ and $\mathbf{V}_{2}^{\star}$ are identifiable. Hence, without loss of generality, we assume $\mathbf{V}_{M}^{\star}=\mathbf{U}_{M}^{\star}\bm{\Lambda}_{M}^{\star}$ for $M\in\{1,2\}$, where $\mathbf{U}_{M}^{\star}\in\mathcal{O}_{d_{M},p}$ contains the left singular vectors and $\bm{\Lambda}_{M}^{\star}\in\mathbb{R}^{p\times p}$ is diagonal. Model (2.1) is also invariant under the shift $\mathbf{v}_{w}^{\star}\leftarrow\mathbf{v}_{w}^{\star}-(1/d_{M})\sum_{w^{\prime}\in{\cal W}^{(M)}}\mathbf{v}_{w^{\prime}}^{\star}$ for all $w\in{\cal W}^{(M)}$, and ${\bm{\Sigma}}_{M}\leftarrow(\mathbf{I}-(1/d_{M})\bm{1}_{d_{M}}\bm{1}_{d_{M}}^{\scriptscriptstyle\sf T}){\bm{\Sigma}}_{M}$. Hence, we may also assume $\mathbf{V}_{M}^{\star{\scriptscriptstyle\sf T}}\bm{1}_{d_{M}}=\bm{0}$ and ${\bm{\Sigma}}_{M}^{\scriptscriptstyle\sf T}\bm{1}_{d_{M}}=\bm{0}$.
@@
Assumption 3.1 is widely known as the incoherence constant condition in the literature appearing in matrix completion (Candès and Recht, , 2009), PCA (Zhang et al., , 2022), and the analysis of representation learning algorithm (Ji et al., , 2023).
@@
Assume that $d_{1}\wedge d_{2}\gg p\log^{2}(n+d)$ and $(d_{1}\wedge d_{2})^{2}\gg p^{4}(d_{1}\vee d_{2})\log(n+d)$, $n\gg p^{2}(d_{1}^{2}\vee d_{2}^{2})\log^{2}(n+d)$, and $n(S_{1}^{(1)}\wedge S_{1}^{(2)})\gg p^{2}(d_{1}^{6}\vee d_{2}^{6})\log^{2}(n+d)$.
Assumption 3.2 imposes a condition on the regime of unique features $d_{1}$ and $d_{2}$, and the number of patients $n$. We will discuss the sufficiency of the condition in Remark 3.1. This assumption is technically necessary in bounding the normalizing constant of the model (Lemma S5.2). We introduce the notation ${\bm{\Sigma}}=\mathrm{diag}({\bm{\Sigma}}_{1},{\bm{\Sigma}}_{2})$.
@@
Assumption 3.3 implies that the signal-to-noise ratio is bounded above and below. Note that the signal, measured by the smallest positive singular value of $(1/p)\mathbf{V}_{M}^{\star}\mathbf{V}_{M^{\prime}}^{\star{\scriptscriptstyle\sf T}}$ for $M,M^{\prime}\in\{1,2\}$ is of order $1/p$ from Assumption 3.1.
@@
$\Big{\|}\mathbb{PMI}^{(M,M)}-\quantity(\frac{1}{p}\mathbf{V}_{M}^{\star}\mathbf{V}_{M}^{\star{\scriptscriptstyle\sf T}}+{\bm{\Sigma}}_{M})\Big{\|}_{\mathrm{F}}$ $\lesssim\frac{p^{2}}{d_{M}}\log(n+d),\ \ M\in\{1,2\}$
$\Big{\|}\mathbb{PMI}^{(1,2)}-\frac{1}{p}\mathbf{V}_{1}^{\star}\mathbf{V}_{2}^{\star{\scriptscriptstyle\sf T}}\Big{\|}_{\mathrm{F}}$ $\lesssim\frac{p^{2}(d_{1}\vee d_{2})}{(d_{1}\wedge d_{2})^{2}}\log(n+d),\quad\text{ and hence }$
$\Big{\|}\mathbb{PMI}-\quantity(\frac{1}{p}\mathbf{V}^{\star}\mathbf{V}^{\star{\scriptscriptstyle\sf T}}+{\bm{\Sigma}})\Big{\|}_{\mathrm{F}}$ $\lesssim\frac{p^{2}(d_{1}\vee d_{2})}{(d_{1}\wedge d_{2})^{2}}\log(n+d).$
The proof of Theorem 3.1 is presented in Supplementary S4.1. It shows that the PMI matrix for each modality can be closely approximated by the product of embeddings plus the noise covariance matrix ${\bm{\Sigma}}_{M}$. When the noise level ${\bm{\Sigma}}_{M}$ is substantial, the PMI value $\mathbb{PMI}^{(M,M)}(w,w^{\prime})$ for a single modality may significantly deviate from $\langle\mathbf{v}_{w}^{\star},\mathbf{v}_{w^{\prime}}^{\star}\rangle/p$. On the other hand, the cross-modal PMI matrix $\mathbb{PMI}^{(1,2)}$ can be directly approximated by $\mathbf{V}_{1}^{\star}\mathbf{V}_{2}^{\star{\scriptscriptstyle\sf T}}/p$. The next theorem affirms that $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$ is an effective estimator for $\mathbb{PMI}^{(1,2)}$.
Assumption 3.4.
For $M=1,2$, assume that $S_{2}^{(M)2}\gg S_{1}^{(M)}$, $S_{1}^{(M)}\gtrsim S_{2}^{(M)}\gtrsim S_{4}^{(M)}$ and $S_{2}^{(1,2)}\gtrsim S_{4}^{(1,2)}$. Also assume the non-negative correlation between $(T_{i}^{(1)})_{i\in[n]}$ and $(T_{i}^{(2)})_{i\in[n]}$, i.e., $S_{2}^{(1,2)2}\geq S_{1}^{(1)}S_{1}^{(2)}$.
Assumption 3.4 requires that the empirical variance of $T_{i}^{(M)}$ dominates the mean of $T_{i}^{(M)}$, and their moments are comparable to each other. For the same length setting $T_{i}^{(M)}\equiv T^{(M)}$ for all $i$, the assumption is satisfied if $T^{(M)}\to\infty$. Assumption 3.4 is satisfied for exponential distribution and gamma distribution with constant order shape parameter.
@@
Under Assumptions 3.1-3.3 and 3.4, with probability $1-\exp(-\Omega(\log^{2}(n+d)))$,
@@
$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}-\mathbb{PMI}^{(1,2)}}_{\mathrm{F}}$ $\lesssim\left\{\frac{d_{1}^{3/2}d_{2}^{3/2}}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{d_{1}^{1/2}d_{2}^{1/2}}{\sqrt{n}}+\frac{d_{1}^{1/2}d_{2}^{1/2}p^{2}}{(d_{1}\wedge d_{2})^{2}}\right\}\log(n+d).$